Jennifer is 6 years older than Sue. In 4 years, she will be twice as old as Sue was 5 years ago. Find their ages now.
Answer by edjones(7311)

Let j be jennifer's age and s be sue's in the present.
A) j=s+6
B) j+4=2(s-5)
.
B) s+6+4=2s-10 Replace j with s+6
s+10=2s-10
20=s add -s and 10 to each side.
j=26 years old.
.
Ed